MILLIMETER WAVE SPECTROSCOPY OF THE VINYL-D (H_{2}CCD) RADICAL GENERATED BY UV LASER PHOTOLYSIS IN A PULSED SUPERSONIC JET EXPANSION. DETERMINATION OF THE PRO-TON TUNNELING SPLITTING

MASATO HAYASHI, KENSUKE HARADA, \underline{KEIICHI TANAKA}, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan; RICHARD LAVRICH, National Institute of Standards and Technology, USA.

The rotational and proton tunneling spectra of the vinyl-d (H_{2}CCD) radical have been observed by millimeter wave spectroscopy combined with a pulsed supersonic jet technique. H_{2}CCD was generated by the 193 nm excimer laser photolysis of vinyl chloride-d(H_{2}CCDCl). The pure rotational transitions (a-type transitions), N_{K_{a}K_{c}}=1_{01}−0_{00}, 2_{02}−1_{01}, 3_{03}−2_{02}, 2_{11}−1_{10}, 2_{12}−1_{11} in the 0^{+} and 0^{−} states, were observed in the frequency region of 52.9−158.7 GHz. The 0^{+} ← 0^{−} and 0^{−} ← 0^{+} proton tunneling transitions (b-type transitions), N_{K_{a}K_{c}}=1_{11}−0_{00}, 1_{10}−1_{01}, 2_{11}−2_{02}, 2_{12}−2_{01}, were observed in the frequency region of 184.7−236.8 GHz. The observed rotational lines were split into fine and hyperfine components due to the spin-rotation interaction and the spin- nuclear spin interaction of the acetylenic (CD) deuteron (α-deuteron) and the methylenic (CH_{2}) protons (β- protons). The moleculer constants such as rotational constants, spin- rotation interaction constants, and hyperfine interection constants, were determined by a least squares fitting of the observed spectrum, together with the proton tunneling splitting ΔE_{0} = 1164.861(20) MHz. The barrier height of the double minimum potential was estimated to be 1520 cm^{−1} from the observed tunneling splitting assuming a one dimensional model. This potential barrier height is consistent with that of the normal species(H_{2}CCH), 1580 cm^{−1}, derived from the tunneling splitting reported to be 16271.8429(59) MHz ^{a}, if the zero point energy is taken into consideration.

aK. Tanaka, M. Toshimitsu, K. Harada, T. Tanaka, J. Chem. Phys. 120, 3604 (2004)